OpenStudy (anonymous):

Find the inverse Laplace transform of \(\frac{1}{s(s^2 + \omega ^2)}\)

OpenStudy (anonymous):

\[L(1*\frac{sin\omega t}{\omega }) = \frac{1}{s(s^2+\omega ^2)}\] But then, I don't know how to continue...

hartnn (hartnn):

i would rather do partial fractions. \(\huge \frac{1}{s(s^2 + \omega ^2)}=\frac{Ax+B}{(s^2 + \omega ^2)}+\frac{C}{s}\)

hartnn (hartnn):

find A,B,C

OpenStudy (anonymous):

Ax+B ?! As+B?!

hartnn (hartnn):

yes, As+B

hartnn (hartnn):

i am getting C=1/w^2 , B=0 , A =-1/w^2 are u getting the same ?

OpenStudy (anonymous):

\[(As+B)s + C(s^2 + \omega^2) = 1\] \[A+C = 0\]\[B =0\]\[C = \frac{1}{\omega^2}\] \[L^{-1}(\frac{1}{s(s^2 + \omega ^2)})=L^{-1}(\frac{-1}{\omega^2(s^2 + \omega ^2)}+\frac{1}{\omega^2s})\]Hmm.. How to do the inverse of Laplace tramsform of \(\frac{-1}{\omega^2(s^2 + \omega ^2)}\) ??

hartnn (hartnn):

did u forget s ?

hartnn (hartnn):

As+B

hartnn (hartnn):

\(L^{-1}(\frac{1}{s(s^2 + \omega ^2)})=L^{-1}(\frac{-s}{\omega^2(s^2 + \omega ^2)}+\frac{1}{\omega^2s})\)

hartnn (hartnn):

w^2 is constant

OpenStudy (anonymous):

Oh!!! Sorry!!!

OpenStudy (anonymous):

\[L^{-1}(\frac{1}{s(s^2 + \omega ^2)})=L^{-1}(\frac{-s}{\omega^2(s^2 + \omega ^2)}+\frac{1}{\omega^2s})=\frac{1}{\omega^2}-\frac{cos\omega t}{\omega^2} = \frac{1-cos\omega t}{\omega^2}\]No wonder why I couldn't get the answer!!! Thanks!! May I know if how to get the answer using the convolution one?

hartnn (hartnn):

i was trying exactly that...using convolution

hartnn (hartnn):

\(\int_0^t u(t-u)sin (\omega u)/\omega \:\:du\) did u reach here ?

hartnn (hartnn):

and i am getting same answer with convolution also.

hartnn (hartnn):

also its , \(\huge L(u(t)*\frac{sin\omega t}{\omega }) = \frac{1}{s(s^2+\omega ^2)}\)

hartnn (hartnn):

if and after u get this : \(\int \limits_0^t u(t-u) \frac{\sin (\omega u)}{\omega} \:\:du\) u just put t-u = something and solve the integral, yo'll get the same answer.....

hartnn (hartnn):

doubts ? i assume you are trying...

OpenStudy (anonymous):

Why is it \( L(u(t)*\frac{sin\omega t}{\omega }) = \frac{1}{s(s^2+\omega ^2)}\) ???

hartnn (hartnn):

because L[u(t)] = 1/s and NOT L[1] = 1/s ...

OpenStudy (anonymous):

!!!!! How come...

hartnn (hartnn):

do u want me to prove that ? i suggest you try first...using definition of Laplace, with u(t), u get 1/s with 1, you won't get 1/s..

OpenStudy (anonymous):

Of course I try first :) Sorry to keep you waiting again!

hartnn (hartnn):

in the definition its 0 to infinity or -infinity to infinity ??

OpenStudy (anonymous):

0 to infinity. That's what I've learnt..

hartnn (hartnn):

okk....then u get same answer for 1 and u(t) because when t>=0, u(t) = 1 then u'll get same answer for both u(t)*... and 1*....

hartnn (hartnn):

so suit yourself...i have habit of taking u(t) because for some transforms, the definition is -infinity to infinity where u(t)and 1 are not same...

OpenStudy (anonymous):

I haven't learnt \(L (u(t)) = \frac{1}{s}\)....

hartnn (hartnn):

then take it as 1...

hartnn (hartnn):

\(\int_0^t 1.sin (\omega (t-u))/\omega \:\:du\) u got this?

OpenStudy (anonymous):

I... think I got that..

hartnn (hartnn):

then its just a normal definite integration problem

hartnn (hartnn):

didn't get it ?

OpenStudy (anonymous):

\[f*g =\int_0^tf(\tau) g(t-\tau) d\tau\] \[\int_0^t 1\cdot \frac{sin (\omega (t-u))}{\omega} du\]\[=\frac{1}{\omega^2}cos (\omega t-\omega u)|_0^t\]\[=\frac{1}{\omega^2}[cos(0) - cos\omega t]=\frac{1-cos\omega t}{\omega^2}\] Wow!!! Thanks a ton!!!!

hartnn (hartnn):

welcome ^_^